OpenStudy (anonymous):

What is the value of the discriminant of the polynomial below? x2 - 9

OpenStudy (anonymous):

I might be wrong but i think the discriminant is the part of the quadratic formula under the sqrt so this would 9^2-4*1*0 because a=1 b=9 and c=0. I might be wrong though.

OpenStudy (anonymous):

nvm I am right

OpenStudy (anonymous):

The discriminant \(\Delta\) of the polynomial \(ax^2+bx+c=0\) comes from the Quadratic formula and is given as follows.\[\Huge \Delta = b^2-4ac\]

OpenStudy (anonymous):

yakeyglee this is what I said.... b^2-4ac is the part under the sqrt of the quadratic formula is it not? and if you plug in the values from the equation above wouldn't it be 9^2-4*1*0 ?

OpenStudy (anonymous):

Yes it is. One could write more simply \(x=\dfrac{-b\pm\sqrt{\Delta}}{2a}\). And your formulation is not quite right. Remember that \(x^2-9\equiv x^2+0x-9\).

OpenStudy (anonymous):

Yeah my bad, but it was obviously a silly mistake. Really? If \[\Delta=b^2-4ac \] that means you can replace b^2-4ac with \[\Delta \] I was not aware of this....

OpenStudy (anonymous):

Well... since you're defining the equality in that manner, you can substitute it.

OpenStudy (anonymous):

Yeah my mind was just blown man =)